OpenStudy (anonymous):

Find the average value of y=sqrt(64-(x^2), on its domain.

OpenStudy (anonymous):

a. 2pi b. 4pi c. 2 d. 4

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

you got the domain?

OpenStudy (anonymous):

-8 to 8

OpenStudy (anonymous):

an interval of length \(16\) right?

OpenStudy (anonymous):

correct @satellite73

OpenStudy (anonymous):

\[\frac{1}{16}\int_{-8}^8\sqrt{64-x^2}dx\] the integral you do by area not by anti derivatives

OpenStudy (anonymous):

2pi or 6.2831...

OpenStudy (anonymous):

right? @satellite73

OpenStudy (anonymous):

the integral is have the area of a circle with radius 8 so \(\frac{1}{2}\times 64\pi=32\pi\)

OpenStudy (anonymous):

yeah divide by \(16\) and get \(2\pi\) what you said

OpenStudy (anonymous):

*half

OpenStudy (anonymous):

yah!! @satellite73 can u help me w/ another problem?

OpenStudy (anonymous):

maybe?

OpenStudy (anonymous):

Find the domain for the particular solution to the differential equation (dy/dx) = (3y/x), with initial condition y(1) = 1.

OpenStudy (anonymous):

a. x > 0 b. x < 0 c. x ≠ 0 d. All real numbers

OpenStudy (anonymous):

i would think the solution is \[y=Cx^3\]right?

OpenStudy (anonymous):

maybe there is something here i am missing

OpenStudy (anonymous):